# The impact of COVID-19 on building energetics Article Type: Brief Communication

### v1

@@ Abstract
Buildings are responsible for a significant portion of global energy demand and GHG emissions. Using the Massachusetts Institute of Technology campus as a case study, we find that, similar to the baseline metabolism of biological organisms, large buildings are on average more energetically efficient per unit size than smaller buildings. This suggests that institutions can be perceived as super populations with buildings as units (organisms) following standard metabolic relationships. Importantly, the relative efficiency of larger buildings progressively increased to until 2020. However, the COVID-19 pandemic acted as a major shock, disrupting this trend and leading to a reversal to the expected 25% baseline level. This suggests that energetic adaptations are contingent on relatively stable conditions.
@@ Introduction
Efforts to address global warming have primarily focused on renewable energy innovation and carbon pricing mechanisms [1, 2], while reducing building energy use has received less attention. Yet, the building sector plays a crucial role in global energy use and carbon emissions, accounting for 21% of global GHG emissions (12 -eq) and 31% of global final energy demand (128.8 EJ) in 2019 [3]. Notably, the global floor area of buildings in cities is projected to increase from 244 billion m2 in 2020 to 427 billion m2 by 2050 [4]. This means that the world is expected to add about 183 billion m2 of new floor area to the global building stock, the equivalent of adding two entire Boston cities every month for 30 years. Therefore, it becomes imperative to understand the physical and socio-economic factors driving building energetics for a sustainable future.
@@
The study of energetics has a long tradition in the life sciences [5, 6, 7]. Evidence shows that biological organisms tend to be more energetically efficient per unit sizeβknown as Kleiberβs Law [8]. Typically, the dependence of metabolism () on body mass () takes the nonlinear form of a power law , where is a taxon-specific normalization constant and represents the scaling exponent [9]. For example, in the animal kingdom, the scaling exponent is typically observed to be sublinear around [10]. This implies that every unit increase in size, there is an average decrease in energy requirements per unit size. Put simply, while an elephant requires more energy than a mouse; gram by gram, the elephant requires less energy than the mouse. Interestingly, the scaling exponent seems to be different across the tree of life. It has been shown that ordering organisms from relatively simple (e.g., prokaryotes) to relatively complex (e.g., metazoans), the scaling exponent changes from superlinear () to sublinear () [10]. These scaling changes are expected to arise due to adaptation processes driven by body plan innovations filtered by natural selection [10, 11, 9].
@@
The scaling patterns observed in biological organisms have also been observed in the energetic requirements of cities [12, 13]. Requirements linked to infrastructure tend to depend sublinearly on citiesβ area, and in contrast requirements linked to socio-economic factors tend to depend superlinearly on citiesβ area [12]. Importantly, it has been shown that such scaling exponents in cities can change relatively fast [14], reflecting the differences in speed between biological and socio-economic adaptations [15]. However, at lower levels of organization where energetic interventions can be more manageable, such as buildings, this analysis has remained more elusive due to the difficulties in collecting fine-grained data [16]. Moreover, it remains unclear the potential impact of strong activity perturbations on building energetics. To address this gap, we use the maintenance energy usage in buildings at the Massachusetts Institute of Technology (MIT) from 2009 to 2024 to analyze energetic trends at the scale of individual buildings[17]. This usage measurement entails the total of both electricity and gas, no information was available for energy use through materials. In particular, we investigate the relationship between energy use and building size, the dynamics of the scaling exponents, and the impact of activity disruptions caused by COVID-19. Lastly, we discuss the implications of our work for sustainable city design.
@@ Results
Focusing our analysis on the energetics of 86 buildings reported in 2024, we found that energy use scales with volume as a power law with scaling exponent of (Fig. 2a) and constant (GJ per unit volume) (Fig. 2b). This reveals that, similar to the baseline metabolism of biological organisms, large buildings are on average more energetically efficient per unit size than small buildings. Because it has become debatable how to better measure the efficiency of buildings [16], this scaling pattern naturally establishes a baseline efficiency for buildings, where deviations from the mean (i.e., positive or negative regression residuals) would imply a more or less efficient building than the baseline according to volume (Fig. 1a shows standardized residuals).
@@
From 2009 to 2024, MIT (hosting almost 30,000 people [17]) had an average energy use in buildings of PJ (a maximum of 2.88 PJ in 2011 and a minimum of 2.5 PJ in 2024). This usage is equivalent to 69,000 average U.S. households. However, changes in energy use across years were not proportional for every building. Figure 2a shows that the scaling exponent progressively decreased from in 2009 to in 2020 (Fig. 2b). In other words, the relative efficiency of large buildings increase to (from the baseline). In contrast, the normalization constant progressively increased from in 2009 to in 2020. That is, the demand for energy per unit volume increased, reflecting a proportionally larger usage of smaller buildings. In sum, these results show that efficiency improvements exhibited by MIT until 2020, were coming precisely from the most energy-demanding buildings.
@@
However, the COVID-19 pandemic led to a reversal to the mean from 2021-2024, bringing efficiency back to the baseline of relative efficiency (Fig. 2c-d). The reversal to the mean effect implies that values far from the expectation in the present tend to be followed by values closer to the expectation in the future [20, 21]. To test this effect, we calculated the correlation between residuals in year 2009 () and subsequent residuals (). This correlation implies the linear relationship , where represents the vector of standardized residuals (Fig. 1a) and corresponds to the Pearson correlation. The weaker the correlation (and statistically detectable), the stronger the reversal to the mean [20]. In line with this concept, Figure 2c shows that correlations were weaker from 2021-2023. Additionally, the reversal to the mean effect requires that the conditional expected value of observations in the future are bounded by the conditional expected value of observations in the present [21]. Using again standardized residuals as observations, the previous statement formally implies and , where is a standardized threshold value (interpreted as standard deviations). Figure 2d shows the strongest reversal to the mean () within the period 2021-2024 across different threshold values.
@@ Discussion
The findings of this study reveal a distinct scaling relationship between building energy use and volume, mirroring patterns observed in biological metabolism [12, 9]. This discovery aligns with well-established metabolic scaling laws, where larger organisms exhibit greater efficiency in energy usage per unit size (Kleiberβs Law [8]). Specifically, the energy use of buildings follows a sublinear power law, indicating that larger buildings are inherently more energy-efficient per unit volume compared to smaller ones. While we did not have information about people in each building (due to the complexity in the day-to-day dynamics at MIT), it is expected that the number of people will grow proportionally to building size. Notably, the decrease in energy use per unit volume in larger buildings likely stems from structural and operational factors, such as centralized circulation systems (less energy dissipation) and economies of scale in energy distribution. Notably, our results reveal that MIT can be characterized as a super-population with buildings as units (organisms) following physico-chemical laws.
Over the years preceding 2020, the relative efficiency advantage of large buildings progressively increased, reflecting an adaptation likely influenced by advancements in energy management and sustainability practices. However, the disruptions caused by COVID-19 abruptly altered this trajectory, leading to a reversal to baseline 25% efficiency levels (Fig. 2b). This alteration may be the consequence of eliminating adaptive practices reducing energy waste over the years. The rapid shift in energy efficiency trends highlights the sensitivity of building energetics to external shocks. This reversal to the mean suggests that, while efficiency gains can be achieved through sustainability-driven targets, they remain contingent on stable usage patterns and adaptive management strategies [22]. In general, this resonates with biological innovations (or adaptations) that tend to occur during periods of relatively stable environments [11, 10]. Yet, under strong perturbations, energetic processes may reverse to the expectations set by first principles.
@@
Acknowledgments. SS acknowledges support from the National Science Foundation under Grant No. DEB-2436069 and MIT Google Program for Computing Innovation. SB acknowlesges support from MIT-IBM Watson AI Lab, MIT J-WAFS, the NSF/NSERC ABC Global Center on AI and Biodiversity Change, and the Schmidt AI2050 Fellowship. We are also grateful to Yu Cheng for guidance on data, and Timothy Gutowsky, Anette (Peko) Hosoi, Julie Newman, Washington Taylor, Maria Yang, and Vicky Yang for insightful discussion that led to the development of this work.

### v2

# The impact of strong activity disruption on building energetics

@@ Abstract
Evidence shows that biological organisms tend to be more energetically efficient per unit size. These scaling patterns observed in biological organisms have also been observed in the energetic requirements of cities. However, at lower levels of organization where energetic interventions can be more manageable, such as buildings, this analysis has remained more elusive due to the difficulties in collecting fine-grained data. Here, we use the maintenance energy usage in buildings at the Massachusetts Institute of Technology (MIT) from 2009 to 2024 to analyze energetic trends at the scale of individual buildings and their sensitivity to strong external perturbations. We find that, similar to the baseline metabolism of biological organisms, large buildings are on average more energetically efficient per unit size than smaller buildings. Because it has become debatable how to better measure the efficiency of buildings, this scaling pattern naturally establishes a baseline efficiency for buildings, where deviations from the mean would imply a more or less efficient building than the baseline according to volume. This relative efficiency progressively increased to until 2020. However, the strong activity disruption caused by the COVID-19 pandemic acted as a major shock, removing this trend and leading to a reversal to the expected 24% baseline level. This suggests that energetic adaptations are contingent on relatively stable conditions.
Keywords: energy, metabolism, scaling, adaptation, society
@@ Introduction
Efforts to address global warming have primarily focused on renewable energy innovation and carbon pricing mechanisms [1, 2], while reducing building energy use has received less attention. Yet, the building sector plays a crucial role in global energy use and carbon emissions, accounting for 21% of global GHG emissions (12 -eq) and 31% of global final energy demand (128.8 EJ) in 2019 [3]. Notably, the global floor area of buildings in cities is projected to increase from 244 billion m2 in 2020 to 427 billion m2 by 2050 [4]. This means that the world is expected to add about 183 billion m2 of new floor area to the global building stock, the equivalent of adding two entire Boston cities every month for 30 years. It becomes imperative then to understand the physical and socio-economic factors driving building energetics for a sustainable future.
@@
The study of energetics has a long tradition in the life sciences [5, 6, 7]. Evidence shows that biological organisms tend to be more energetically efficient per unit size—known as Kleiber’s Law [8]. Typically, the dependence of metabolism () on body mass () takes the nonlinear form , where is a taxon-specific normalization constant and represents the scaling exponent [9]. For example, in the animal kingdom, the scaling exponent is typically observed to be sublinear around [10]. This implies that every unit increase in size, there is an average decrease in energy requirements per unit size. Put simply, while an elephant requires more energy than a mouse; the elephant can be sustained more efficiently than the mouse. Interestingly, the scaling exponent seems to be different across the tree of life. It has been shown that ordering organisms from relatively simple (e.g., prokaryotes) to relatively complex (e.g., metazoans), the scaling exponent changes from superlinear () to sublinear () [10]. These scaling changes are expected to arise due to adaptation processes driven by body plan innovations filtered by natural selection [10, 11, 9].
@@
The scaling patterns observed in biological organisms have also been observed in the energetic requirements of cities [12, 13]. Requirements linked to infrastructure tend to depend sublinearly on cities’ area, and in contrast requirements linked to socio-economic factors tend to depend superlinearly on cities’ area [12]. Importantly, it has been shown that such scaling exponents in cities can change relatively fast [14], reflecting the differences in speed between biological and socio-economic adaptations [15]. However, at lower levels of organization where energetic interventions can be more manageable, such as buildings, this analysis has remained more elusive due to the difficulties in collecting fine-grained data [16]. Indeed, a building as a physical individual object does not undergo significant evolution (though one can argue that modern retrofit and improvements all count as evolution, in fact, a lot of iconic buildings cost more and more for modern improvements nowadays), but a building as a human-made system has gone through human selection since early settlement, from the makeshift tents in Stone Age (million years ago- year ago) to the modern day architecture. Moreover, it remains unclear the potential impact of strong activity disruptions on building energetics. To address this gap, we use the maintenance energy usage in buildings at the Massachusetts Institute of Technology (MIT) from 2009 to 2024 to analyze energetic trends at the scale of individual buildings[17]. This usage measurement entails the total of both electricity and gas, no information was available for energy use through materials. In particular, we investigate the relationship between energy use and building size, the dynamics of the scaling exponents, and the impact of activity disruptions caused by COVID-19. Lastly, we discuss the implications of our work for sustainable city design.
@@ Results
Focusing our analysis on the energetics of 86 buildings reported in 2024, we found that energy use has a scaling relationship with volume defined by an exponent of (Fig. 2a) and constant (GJ per unit volume) (Fig. 2b). This reveals that, similar to the baseline metabolism of biological organisms, large buildings are on average more energetically efficient per unit size than small buildings. Because it has become debatable how to better measure the efficiency of buildings [16], this scaling pattern naturally establishes a baseline efficiency for buildings, where deviations from the mean (i.e., positive or negative regression residuals) would imply a more or less efficient building than the baseline according to volume (Fig. 1a shows standardized residuals).
@@
From 2009 to 2024, MIT (hosting almost 30,000 people [17]) had an average energy use in buildings of PJ (a maximum of 2.88 PJ in 2011 and a minimum of 2.5 PJ in 2024). This usage is equivalent to 69,000 average U.S. households. However, changes in energy use across years were not proportional for every building. Figure 2a shows that the scaling exponent progressively decreased from in 2009 to in 2020 (Fig. 2b). In other words, the relative efficiency of large buildings increase to (from the baseline). In contrast, the normalization constant progressively increased from in 2009 to in 2020. That is, the demand for energy per unit volume increased, reflecting a proportionally larger usage of smaller buildings. Note that these statistical values are typically considered meaningful in the area of biology and social sciences (given the noise or scarcity of the data) [14]. In sum, these results show that efficiency improvements exhibited by MIT until 2020, were coming precisely from the most energy-demanding buildings.
@@
However, the strong activity disruption caused by the COVID-19 pandemic led to a reversal to the mean from 2021-2024, bringing efficiency back to the baseline of relative efficiency (Fig. 3a-b). The reversal to the mean effect implies that values far from the expectation in the present tend to be followed by values closer to the expectation in the future [20, 21]. To test this effect, we calculated the correlation between residuals in year 2009 () and subsequent residuals (). This correlation implies the linear relationship , where (respectively, ) represents the vector of standardized residuals (Fig. 1a) and corresponds to the Pearson correlation. The weaker the correlation (and statistically detectable), the stronger the reversal to the mean [20]. In line with this concept, Figure 3a shows that correlations were weaker from 2021-2023. Additionally, the reversal to the mean effect requires that the conditional expected value of observations in the future are bounded by the conditional expected value of observations in the present [21]. Using again standardized residuals as observations, the previous statement formally implies and , where is a standardized threshold value (interpreted as standard deviations). Figure 3b shows the strongest reversal to the mean () within the period 2021-2024 across different threshold values.
@@ Discussion
The findings of this study reveal a distinct scaling relationship between building energy use and volume, mirroring patterns observed in biological metabolism [12, 9]. This discovery aligns with well-established metabolic scaling relationships, where larger organisms exhibit greater efficiency in energy usage per unit size (Kleiber’s Law [8]). Specifically, the energy use of buildings follows a sublinear scaling as a functio of size, indicating that larger buildings are inherently more energy-efficient per unit volume compared to smaller ones. While we did not have information about people in each building (due to the complexity in the day-to-day dynamics at MIT), it is expected that the number of people will grow proportionally to building size. Notably, the decrease in energy use per unit volume in larger buildings likely stems from structural and operational factors, such as centralized circulation systems (less energy dissipation) and economies of scale in energy distribution. Future work should further investigate the extent to which human-made systems can indeed have similar energetic developments as biological organisms [12].
Over the years preceding 2020, the relative efficiency advantage of large buildings progressively increased, reflecting an adaptation likely influenced by advancements in energy management and sustainability practices. However, the strong activity disruptions caused by the COVID-19 pandemic abruptly altered this trajectory, leading to a reversal to baseline 24% efficiency levels (Fig. 2b). This alteration may be the consequence of eliminating adaptive practices reducing energy waste over the years. The rapid shift in energy efficiency trends highlights the sensitivity of building energetics to external shocks. This reversal to the mean suggests that, while efficiency gains can be achieved through sustainability-driven targets, they remain contingent on stable usage patterns and adaptive management strategies [22]. In general, this resonates with biological innovations (or adaptations) that tend to occur during periods of relatively stable environments [11, 10]. Yet, under strong perturbations, energetic processes may reverse to the expectations set by first principles.
@@
Abbreviations
MIT: Massachusetts Institute of Technology
GHG: Green House Gas
US: United States of America
Declarations
Ethics approval and consent to participate This study does not incorporate any personal information.
Consent for publication. All of the authors consent to publication.
Availability of data and material: The code and data associated with this work can be accessed at https://github.com/hsuanmina/building-energetics. The code and data will be deposited at Zenodo upon acceptance.
Competing interests. The authors declare no competing interests.
Funding. SS acknowledges support from the National Science Foundation under Grant No. DEB-2436069 and MIT Google Program for Computing Innovation. SB acknowlesges support from MIT-IBM Watson AI Lab, MIT J-WAFS, the NSF/NSERC ABC Global Center on AI and Biodiversity Change, and the Schmidt AI2050 Fellowship.
Author contributions: SS designed and supervised the work. YHH performed the work. All authors analyzed results. SS and YHH wrote a first version of the manuscript. All authors contributed to the revision of the manuscript.
Acknowledgments. We are grateful to Yu Cheng for guidance on data, and Timothy Gutowsky, Anette (Peko) Hosoi, Julie Newman, Washington Taylor, Maria Yang, and Vicky Yang for insightful discussion that led to the development of this work.